RRRG new website top logo resized

 

Restoration diary

Reports of restoration work on 50029 and 50030.

Announcement of working party dates and details of restoration work carried out at working parties.

June 26th 2011

Ian & Pete stretched another two pistons' worth of stretch bolts, before having to abandon this until a spanner can be bent, as the power unit internal structural webbing is now getting in the way of using the spanner properly. They then helped Dave loosen the bolts on the spare bogies to allow the traction motors to be removed sometime this year for overhaul. I test fitted the other two small boards that go behind the divert contactors, and found they fit easily. These two boards house the ETH overvolt relay HOVR & it's associated resistor Z24, the ETH voltage sensing resistor Z23, the ETH earth parallel resistor Z69, the ETH earth fault relay HEFR, & it's associated resistor Z62, and the ETH detector relay HDR. I bolted the left-hand board in place and wired in HOVR, Z23, Z24 & Z69.

Tags:
Continue reading
3335 Hits
0 Comments

June 20th-23rd 2011

Chris & Tom spent four days on site doing various jobs including repairing, straightening & painting the mounting pedestal for the load limiting pot, cabside corrosion repairs to No.2 cab of Repulse, collecting the spare auxiliary gen from Carnforth & taking it to a secure location, continuing to see if a new aux contact mechanism can be made for the reverser for Renown, creating shelving for me in the air filter area of Repulse and various other jobs.

Tags:
Continue reading
13011 Hits
0 Comments

June 12th 2011

Dave as seems to be the case most weekends at the moment, spent both days doing various jobs, then helped Ian & Pete stretch the stretch bolts on another three power unit big ends on a miserable Sunday when it rained continuously. I wired the auxiliary contacts on BMR1, GFR & OIR relays and the Z59 resistor on the board behind the fuse cabinet in Repulse.

Tags:
Continue reading
3490 Hits
0 Comments

June 4th-5th 2011

Dave was on site all weekend whilst Steve & Sarah spent Saturday afternoon and all day Sunday there. Dave carried out various work on brake rigging and some bodywork both on Repulse. He also looked into preparing the traction motors for lifting from the two spare bogie sets, I think the fixing bolts are proving very tough!!!


Steve & Sarah continued identifying the cut cables in Renown. The going was very slow this weekend as the cover to the water header tank level switch would not budge so in the end with help from Dave the whole level switch assembly was removed from the tank. Once the whole assembly was removed the cover was prised away to reveal the cable connections. The two level switch cables were identified and labelled. The level switch operation was tested and found to be good and therefore the whole assembly was reinstalled.


A further 2 cables was identified at the Fire Alarm Relay No1. All the cables were retested at the AWS Relay Box to prove we had identified these.
Steve and Sarah took an hour out on Sunday to meet up with John Patrick - Peak Rail Safety Officer to complete their Personal Track Safety certification.

Tags:
Continue reading
3350 Hits
0 Comments

May 21st-25th 2011

Dave, Chris & Tom spent a marathon five days on site. Dave sorted the spare brake rigging and began cleaning & un-sticking various parts, Tom continued working on the interior wood surrounds on the quarter-lights & drop-lights in 30's No.2 cab & Chris fitted the three remaining sections of 185mm cables to the field divert contactors, as he & I had decided a fresh pair of eyes might be able to devise a solution from pictures he had taken of the cables in 50050. He has also taken the auxiliary contacts off the 'proper' 50 reverser, to see if he can fabricate a new mechanism for the second reverser for Renown & begun cleaning debris from the No.1 end clean air compartment.

Tags:
Continue reading
3487 Hits
0 Comments